Grade 12General Physics

A uniform chain of length 2L is hanging in equilibrium position, if end B is given a slightly downward displacement
the imbalance causes an acceleration. Here pulley is small and smooth & string is inextensible
The velocity v of the string when it slips out of the pulley (height of pulley from floor > 2L)
(A) √(gL/2)
(B) √(2gL)
(C) √(gL)
(D) none of these

Navneet Kumar
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 
\lambda(L+x)g-T=\lambda(L+x)a
T-\lambda(L-x)g=\lambda(L-x)a
Where lambda is mass per unit length of both ends 
Solve these equations by adding them give you
2xg=2La
(x/L)g=a
And  these question is specially check you concept in NLM
 
Regards
Arun